OpenStudy (anonymous):

What is the value of 7P5? A.5040 B.2520 C.42 D.21

jimthompson5910 (jim_thompson5910):

n P r = (n!)/((n-r)!) 7 P 5 = (7!)/((7-5)!) 7 P 5 = (7!)/(2!) 7 P 5 = (7*6*5*4*3*2!)/(2!) 7 P 5 = 7*6*5*4*3 7 P 5 = 2520

OpenStudy (anonymous):

thanks not that good with that

jimthompson5910 (jim_thompson5910):

yw